solve the system of equations
x + 3y = 32
-3x + 2y = 3
and solve for this system of equations
x + y = 5
x - y = -9

Answer by sofiyac(983)

post problems one at a time
solve the system of equations
x + 3y = 32
-3x + 2y = 3
************
3x+9y=96
-3x + 2y = 3
***********
11y=99
y=9
x=32-3y
x=32-3*9
x=32-27
x=5
(5,9)
and solve for this system of equations
x + y = 5
x - y = -9
*************
2x=-4
x=-2
y=5-x
y=5+2
y=7
(-2,7)